OpenStudy (anonymous):

Solved the system by substitution.if the system is inconsistent and has no solution,state this .if the system is dependent,write the form of the solution for any real number x. x+y=50 0.25x+0.75y=0.60(50)

OpenStudy (mathstudent55):

Are you familiar with the substitution method?

OpenStudy (anonymous):

I am not thats why I need help

OpenStudy (mathstudent55):

Ok, no problem. The substitution method is the following: 1. Solve one equation for one variable. 2. Substitute what that variable is equal to in the other equation. Now you have one equation in one variable, so you can find the value of one variable. 3. Substitute the value of the known variable in one of the original equations to find the other variable.

OpenStudy (anonymous):

Can you solved this so as I can learn from it

OpenStudy (mathstudent55):

Ok, now let's go through each step and do it together.

OpenStudy (anonymous):

Ok good

OpenStudy (mathstudent55):

1. Let's solve the first equation for x. \(x + y=50\)

OpenStudy (mathstudent55):

We subtract y from both sides to get: \(x = 50 - y\)

OpenStudy (mathstudent55):

That takes care of step 1. 2. Now we substitute what x is equal to in the second equation.

OpenStudy (mathstudent55):

Here is the second equation. We will insert 50 - y for x, and we can also multiply out the right side. \(0.25x+0.75y=0.60(50)\) \(0.25(50 - y) + 0.75y=30\)

OpenStudy (mathstudent55):

Now we have one equation in one variable, so we solve it for y.

OpenStudy (mathstudent55):

\(0.25(50 - y) + 0.75y=30\) Distribute the 0.25: \(12.5 - 0.25y + 0.75y = 30 \) Combine like terms on the left side: \(12.5 + 0.5y = 30 \) Subtract 12.5 from both sides: \(0.5y = 17.5\)

OpenStudy (anonymous):

How did you get 30

OpenStudy (mathstudent55):

Divide both sides by 0.5: \(y = 35\)

OpenStudy (mathstudent55):

Look at the first line that has 30 in it, and then look at the line just above.

OpenStudy (anonymous):

Oh Ok

OpenStudy (anonymous):

is the system inconsistent and has no solution,or dependent

OpenStudy (mathstudent55):

30 is simply 0.60 * 50

OpenStudy (anonymous):

the system is inconsistent and has no solution,state this .if the system is dependent

OpenStudy (mathstudent55):

Now we move on to step 3. Since we know that y = 35, we substitute y with 35 in one of the two original equations. Let's use the first equation: \(x + y = 50\) \(x + 35 = 50\) Subtract 35 from both sides to get: \(x = 15\) The final answer is: the solution is x = 15 and y = 35.

OpenStudy (anonymous):

so its a solution

OpenStudy (mathstudent55):

Yes. This system has one solution. It is not inconsistent, and it is not dependent.

OpenStudy (anonymous):

I have any equation 9x+4y=23 5x-2y=-2

OpenStudy (anonymous):

But let me solved it first

OpenStudy (mathstudent55):

ok

OpenStudy (anonymous):

9x+4y=23 5x-2=-2 5(23-4y)-2=-2 115-20-2=-2 115-22=-2 -22=-2 -22y=-117 =5.31

OpenStudy (mathstudent55):

Not quite. Remember, you must solve one equation for one variable. It looks like you were trying to solve the first equation for x, to then insert in the second equation. Let's do that together.

OpenStudy (anonymous):

Ok good but am trying

OpenStudy (mathstudent55):

Solve for x: 9x+4y=23 First, we subtract 4y from both sides: 9x = 23 - 4y This is what you got. It is correct up to here, but you have not solved the equation for x. You only have 9x so far. You need one more step. You need to divide both sides by 9. \(x = \dfrac{23}{9} - \dfrac{4}{9}y\) Now it is solved for x.

OpenStudy (mathstudent55):

I see you are tryong. It was a good effort. Now we will iron out the problems.

OpenStudy (mathstudent55):

Now that we have one equation solved for x, we insert what x is equal to in the other equation. This is the "substitution" step, and what gives this method the name "substitution method."

OpenStudy (mathstudent55):

x is equal to the quantity in red. \(x = \color{red}{\dfrac{23}{9} - \dfrac{4}{9}y}\) Here is the second equation. \(5\color{red}{x}-2y=-2\) In the next step, we'll replace the red x by the red quantity. \(5 \left(\color{red}{\dfrac{23}{9} - \dfrac{4}{9}y} \right) - 2y = -2\)

OpenStudy (mathstudent55):

Now we need to solve that equation for y.

OpenStudy (anonymous):

How did we solved for Y

OpenStudy (mathstudent55):

We distribute the 5: \( \dfrac{115}{9} - \dfrac{20}{9}y - 2y = -2 \)

OpenStudy (anonymous):

5x23=115

OpenStudy (anonymous):

5x4=20

OpenStudy (mathstudent55):

Since we have a denominator of 9 on the left side, we multiply the entire equation by 9: \( 9\left[\dfrac{115}{9} - \dfrac{20}{9}y - 2y \right]= 9(-2)\) \(115 - 20y - 18y = -18 \) \(115 - 38y = -18\) \(-38y = -133\) \(y = 3.5\)

OpenStudy (mathstudent55):

Now that we have y = 3.5, we substitute that value into the first equation: \(9x + 4y = 23\) \(9x + 4(3.5) = 23 \) \(9x + 14 = 23\) \(9x = 9\) \(x = 1\)

OpenStudy (mathstudent55):

The solution is x = 1 and y = 3.5.

OpenStudy (anonymous):

How did you get 14

OpenStudy (mathstudent55):

4 times 3.5 = 14

OpenStudy (anonymous):

Oh great is the system is inconsistent and has no solution,state this .if the system is dependent

OpenStudy (mathstudent55):

If you get a solution, that means the system is not inconsistent or dependent.

OpenStudy (anonymous):

Ok thanks

OpenStudy (anonymous):

I have two equation but they are in word problem ...same equation

OpenStudy (mathstudent55):

Start a new post for each question.

OpenStudy (anonymous):

Ok